OpenStudy (anonymous):

Solve for x. x^2 - 8x + 7 = 0

OpenStudy (anonymous):

i think i got the answer but im not sure.

OpenStudy (anonymous):

x =-b+-sqrt(b^2-4ac) all over 2a where ax^2+bx+c=0

OpenStudy (anonymous):

is it (x-1)(x-7)?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

x=1 or x=7

OpenStudy (anonymous):

dont i have to like turn them to positive or something?

OpenStudy (anonymous):

so i solve it for the 0 right?

OpenStudy (anonymous):

Yeah alright so this is how it works: you have (x-a)(x-b)=0

OpenStudy (anonymous):

In order for that equation to be true, you have to either have x-a=0 or x-b=0, so set each of them equal to 0 and solve for x This is because when x=a, (a-a)(a-b)=0, 0(a-b)=0, 0=0 (it works) x=b, (b-a)(b-b)=0, (b-a)0=0, 0=0 (it works)

OpenStudy (anonymous):

Hope this helped :)

OpenStudy (anonymous):

lol dont really understand it your way, but thanks anyways. =D

OpenStudy (anonymous):

You just set each of the things in the ( )'s to 0 when you have it in the (x-a)(x-b)=0 :P because when one of the ( ) terms is 0, the expression is true